OpenStudy (anonymous):

2i + (x + yi) = 6, what is x + yi?

OpenStudy (anonymous):

x+yi=complex number

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you wanna find x?

OpenStudy (anonymous):

and yi

OpenStudy (anonymous):

aaah wou wanna find x+yi

OpenStudy (anonymous):

yes. I don't unserstand

OpenStudy (anonymous):

note x+yi=z 2i+z=6 \[(2i=6-z)|^{2}\] note that i^2=-1 \[-4=36-12z+z^2\] \[z^2-12z+32=0\] with the solutions \[z _{1}=4;z _{2}=8\]

OpenStudy (anonymous):

so x+yi can be 4 or 8

OpenStudy (anonymous):

I see thank you!